Circular No. 6365

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) C. G. Mason, R. D. Gehrz, T. J. Jones, and D. Williams, University

of Minnesota (UM), report infrared photometry of comet C/1996 B2 using the 0.76-m telescope (+ UM bolometer) at O’Brien Observatory: “Photometric magnitudes in a 27′′ circular aperture on Mar. 21.4 UT, when the throw between the source and reference beams was 212′′ (and uncertainties, including systematic errors, were < 10 percent): [1.2 μm] = +4.74, [1.6 μm] = +4.67, [2.3 μm] = +3.96, [3.6 μm] = +2.80, [4.9 μm] = +0.07, [7.8 μm] = −3.24, [8.7 μm] = −3.95, [9.8 μm] = −4.58, [10.3 μm] = −4.71, [11.6 μm] = −5.18, [12.5 μm] = −5.29; [18 μm] = −6.18. With an 84′′ throw (and uncertainties < 20 percent), the magnitudes on Mar. 23.3 were: [1.6 μm] = +4.67, [2.3 μm] = +5.13, [3.6 μm] = +2.63, [4.9 μm] = +0.39, [7.8 μm] = −2.57, [8.7 μm] = −3.29, [9.8 μm] = −3.77, [10.3 μm] = −3.81, [11.6 μm] = −4.08, [12.5 μm] = −4.54, [18 μm] = −5.69. Infrared spectral energy distributions for both dates show that the dust temperature is ∼ 30 percent higher than would be expected for small black spheres at the same heliocentric distance. The 10- and 20-μm silicate emission features are very weak on both dates, and the dust albedo is ∼ 0.15. The dispersion in several measurements made on Mar. 23 suggests that the activity of the nucleus might be variable on short timescales.”

NOVA CASSIOPEIAE 1995 T. Iijima and L. Rosino, Asiago Astrophysical Observatory, report:

“The spectral variation of this nova is being monitored at Asiago Observatory. At the maximum brightness in 1995 mid-December, a pure absorption spectrum of an F-type supergiant appeared (e.g., Dec. 19.8 UT). This absorption spectrum lasted a relatively short time, and spectra taken on 1996 Jan. 2.9 again showed the prominent emission lines of H I, He I, Fe II, etc., which spectroscopically confirmed that this object is not a symbiotic nova but rather a classical slow nova. The blueshift of the P-Cyg-type absorption component and the FWHM of the prominent emission lines were typically −360 ± 30 km/s and 270 ± 20 km/s in the pre-maximum stage (e.g., 1995 Oct. 17.1). These quantities increased to −1060 ± 30 km/s and 640 ± 30 km/s, respectively, in the post-maximum stage (e.g., 1996 Mar. 1.8). This nova is still in the pre-nebular stage in the middle of March 1996.”

1996 March 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green